1. Stock A has a beta of 2.6 and an expected return of 13.6 percent. Stock B has a beta of 1.18 and an expected return of 16.70 percent. At what risk-free rate would these two stocks be correctly priced?

19.28 percent

17.98 percent

17.10 percent

18.67 percent
Answer #1

A. 19.28%

(0.136 – rf) / 2.6 = (0.167 – rf) / 1.18

1.18(0.136 – rf) = 2.6(0.167 – rf)

0.16048 - 1.18rf = 0.4342 - 2.6rf

2.6rf - 1.18rf = 0.4342 - 0.16048

1.42rf = 0.27372

rf = 0.1928 or 19.28%